# TimeFlow: Longitudinal Brain Image Registration and Aging Progression Analysis

### v2

@@ Abstract
Predicting future brain states is crucial for understanding healthy aging and neurodegenerative diseases. Longitudinal brain MRI registration, a cornerstone for such analyses, has long been limited by its inability to forecast future developments, reliance on extensive dense longitudinal data, and the need to balance registration accuracy with temporal smoothness. In this work, we present TimeFlow, a novel framework for longitudinal brain MRI registration that overcomes all these challenges. TimeFlow leverages a U-Net architecture with temporal conditioning inspired by diffusion models, enabling accurate registration using only two images as input and facilitating prospective analyses through future image prediction. Unlike traditional methods, TimeFlow eliminates the demand for explicit smoothness regularizers and dense sequential data while maintaining temporal consistency and continuity. Experimental results highlight its superior performance in both future timepoint prediction and registration accuracy compared to state-of-the-art methods. Additionally, TimeFlow supports novel biological brain aging analyses, effectively differentiating neurodegenerative conditions from healthy aging, all without requiring segmentation thus avoiding non-trivial annotation and inconsistent segmentation flaws. This framework paves the way for accurate, data-efficient, and annotation-free prospective analyses of brain aging and chronic diseases.
Figure 1: Outstanding performance of TimeFlow for future brain MRI prediction. TimeFlow demonstrates the superior capability of future extrapolation compared to (i) NODER, which employs NeuralODE for geodesic regression, and (ii) VoxelMorph with linear scaling. Prediction errors in these methods are marked with red arrows. Given two previous timepoints (Baseline and Year 3), TimeFlow can extrapolate precise and smooth deformation to predict brain MRI in Year 6.
@@ 1 Introduction
Understanding how the human brain changes over time — both in the course of healthy aging and under the influence of neurodegenerative diseases — remains a central challenge in neurology and related fields [33, 34]. Such insights not only shed light on the underlying pathological processes but also pave the way for more individualized treatment planning and earlier, more targeted interventions [6]. Magnetic Resonance (MR) imaging has become the standard modality for capturing these longitudinal changes, owing to its high spatial resolution and excellent soft-tissue contrast. By repeatedly scanning the same individual, clinicians and researchers can visualize evolving patterns of atrophy, lesion progression, and other structural transformations as they unfold [21, 2].
@@
Despite advances in longitudinal registration [17, 43], which aligns a subject’s brain images across multiple time points, two critical challenges persist: 1. Predicting future brain states and analyzing aging progression is one of the ultimate goals of longitudinal registration. While retrospective assessments of past changes are valuable, the ability to predict future brain transformations would dramatically enhance clinical utility [27] and enable clinicians to forecast atrophy trajectories and guide more effective interventions [23]. By extrapolating physically plausible transformations, longitudinal registration enables future scan prediction through simple image warping, avoiding the need for complex generative approaches. Moreover, comparing the predicted future brain changes with the actual observed brain state allows clinicians to assess whether the disease is advancing at an accelerated rate compared to historical progression. This insight can provide early warnings of disease exacerbation and help refine patient-specific prognostic models. Recent works [50, 3, 51] primarily focus on modeling the observed sequence of scans, allowing interpolation between observed timepoints but lacking the capability to extrapolate beyond the final scan, as illustrated in Fig. 1. This limitation makes it difficult to assess deviations from expected progression, anticipate disease worsening, and adjust treatment accordingly. 2. Ensuring temporal continuity with limited sequential data. Existing methods impose temporal smoothness constraints [49, 53, 9] and often require densely sampled longitudinal data (often more than five scans) [28, 36], which is impractical in typical clinical scenarios where only a few follow-up scans (<3) are available.
We introduce TimeFlow, a novel longitudinal brain MRI registration capable of predicting future brain changes. The key contribution is a novel constraint that ensures temporal consistency and continuity — removing the demand for smoothness regularizers. Moreover, TimeFlow can operate longitudinal registration with only paired images in inference, eliminating the requirement for long image sequences. Finally, TimeFlow can extrapolate to future time points and analyze aging progression without relying on segmentation masks. Our contributions are twofold:
@@
    Inspired by conditioning mechanisms used in diffusion models, TimeFlow employs a U-Net architecture conditioned on a temporal parameter reflecting biological aging. This temporal conditioning allows for generating temporally continuous and smooth deformation fields, even when only two images are used as input. Moreover, this approach enables future brain images prediction, surpassing state-of-the-art methods in accuracy and smoothness.
@@
    We develop a novel methodology for longitudinal brain aging and disease progression analysis, utilizing TimeFlow’s advanced predictive capabilities. Our approach quantifies relative aging progression using healthy subjects as a reference baseline, where biological aging is assumed to align with chronological aging. This method enables the identification and differentiation of subjects with Mild Cognitive Impairment (MCI) and Dementia, who typically exhibit accelerated biological aging compared to chronological aging. Remarkably, our method performs this analysis solely through registration, eliminating the need for brain structure annotations and avoiding potential segmentation-related error propagation.
-sho
@@ 2 Related Work
Cross-sectional registration
@@
methods laid the groundwork for medical image registration, pioneering this field with techniques such as parametric spline-based models [41, 29], optical flow [35, 52], and diffeomorphic transformations [44, 47, 5]. However, they demand prohibitive computing time in the order of several hours to register one single subject, posing a significant barrier to clinical deployment. To address this, learning-based frameworks [4, 31, 37, 7, 14, 22] have emerged, dramatically reducing inference time and often improving registration accuracy. Nevertheless, all such methods rely on carefully tuned smoothness regularizers to ensure anatomically plausible deformation fields. An improper weighting factor of smoothness regularizers can either reduce fidelity or introduce unrealistic warps. Additional tuning or adaptive strategies are often needed to strike an optimal balance [18, 32, 48].
Longitudinal registration
@@
methods [42, 19, 20, 11, 12, 36, 9, 24] often leverage the backbone architecture of the cross-sectional registration, whereas one additional constraint is introduced: the temporal continuity/consistency across multiple timepoints of a subject. Two major strategies are often used to tackle this problem. First, an explicit smoothness constraint along the temporal dimension is incorporated to maintain coherent deformation fields over time [53, 28]. However, as with spatial smoothness, temporal regularization faces a similar trade-off: enforcing temporal continuity can inadvertently lower spatial accuracy. Second, many longitudinal methods require multiple, densely sampled scans (>10) to capture temporal dynamics [3, 50]. In real-world clinical scenarios, patients often have only a few follow-up scans ($\leq 3$), restricting model flexibility and practical implementation. Importantly, these methods commonly can only carry out retrospective analyses and are not capable of prospective studies such as future image extrapolation and aging/disease progression risk prediction.
@@ 3 Methods
Figure 2: Left: Architecture of TimeFlow. TimeFlow $f_{\theta}$ integrates time $t$ information via adaptive instance normalization. The network takes the time $t$, the images at first $I_{0}$ and last visit $I_{L}$ as input, and predicts the temporal deformations. Examples of $t=0.5$ (interpolation at middle), $t=1$ (endpoint registration), and $t=2$ (extrapolation) are shown. Right: Interpolation (upper yellow) and Extrapolation (bottom green) Consistency. TimeFlow can predict bidirectional time-continuous deformation fields. The symmetric midpoint similarity forces the deformed images to be aligned at time $t_{k}$, while the midpoint interpolation flow loss encourages the deformation composition consistency. The extrapolation similarity and flow enforcement give information on what a realistic and reasonable extrapolation should be. Training with both consistency losses enables TimeFlow to generate plausible deformation in $t\in[-\infty,+\infty]$. The input images are highlighted by blue squares (Input), while the supervision are highlighted by dot-lined squares (Sup.).
@@ 3.1 Time-Continuous Registration
Given a sequence of longitudinal images $\{I_{k}\mid k=0,\ldots,L\}$, we aim to find a temporally continuous deformation trajectory that connects all the discrete time points. To achieve this, we construct a time-continuous registration function $\phi_{t}=f_{\theta}(I_{0},I_{L},t)$. It takes the first visit $I_{0}$ and the last visit $I_{L}$ as input images and predicts the deformation $\phi_{t}$ at time $t$. As shown in Figure 2, we utilize a time-embedded UNet, frequently used in diffusion models [15, 40] to learn the mapping function. The time variable $t$ is first projected to the latent embedding space by a sinusoidal positional encoding layer and two-layer MLPs. Next, the time latent vector is fed to two separate MLPs to steer the mean and variance of the inter-layer adaptive instance normalization. The network is thus instilled with the capability of generating meaningful time embedding and predicting time-continuous flow fields by the novel time conditioning mechanism. Moreover, to ensure the predicted deformation trajectory passes through each visit $I_{k}$ at time $t=t_{k}$, we incorporate the intermediate time points as supervision in training.
3.2 Interpolation Constraints
Intermediate Similarity.
Our goal is to learn a time-continuous registration function $\phi_{t}^{0\rightarrow L}=f_{\theta}(I_{0},I_{L},t)$. When $t=0$, we have $f_{\theta}(I_{0},I_{L},0)=\textrm{Id}$, the identity mapping. When $t\in(0,1]$, $\phi_{t}^{0\rightarrow L}$ is the forward deformation mapping from $I_{0}$ towards $I_{L}$, and $\phi_{1}^{0\rightarrow L}$ moves $I_{0}$ to reach $I_{L}$. Similarly, for negative values, $t<0$, $\phi_{t}$ maps backward from $I_{L}$ to $I_{0}$. Exploiting bidirectional time-continuous deformations, we can formulate a symmetric similarity measure:
@@ 3.2 Interpolation Constraints
which deforms both images to remove the bias introduced by single-side warping [1, 30, 16]. As shown in the yellow part in Fig. 2(b), $I_{0}\circ\phi_{t_{k}}^{0\rightarrow L}$ and $I_{L}\circ\phi_{t_{k}-1}^{0\rightarrow L}$ describe the deformation of both $I_{0}$ and $I_{L}$ to the time $t=t_{k}$. Given an acquired MRI $I_{k}$ at time $t=t_{k}$, we can simply use it to compute the symmetric intermediate similarity and thus force the network to learn plausible deformation trajectories from $I_{0}$ to $I_{L}$ that pass through intermediate images.
Generalized Intermediate Similarity.
However, in real-world clinical scenarios, many subjects will only have MRI from two time points available. In such a case, we rewrite the symmetric similarity as:
@@
where $\hat{t}$ is sampled from a uniform distribution $\mathcal{U}(0,1)$. The registration is guided properly through this synthetic variant of the symmetric similarity, even without any intermediate observation as supervision. Moreover, taking advantage of the freedom to choose $\hat{t}$, the network can learn densely on the time axis and is no longer constrained to intermediate visits at specific times in the dataset.

### v3

# IEEEexample:BSTcontrol

@@ Abstract
Longitudinal brain analysis is essential for understanding healthy aging and identifying pathological deviations. Longitudinal registration of sequential brain MRI underpins such analyses. However, existing methods are limited by reliance on densely sampled time series, a trade-off between accuracy and temporal smoothness, and an inability to prospectively forecast future brain states. To overcome these challenges, we introduce TimeFlow, a learning-based framework for longitudinal brain MRI registration. TimeFlow uses a U-Net backbone with temporal conditioning to model neuroanatomy as a continuous function of age. Given only two scans from an individual, TimeFlow estimates accurate and temporally coherent deformation fields, enabling non-linear extrapolation to predict future brain states. This is achieved by our proposed inter-/extrapolation consistency constraints applied to both the deformation fields and deformed images. Remarkably, these constraints preserve temporal consistency and continuity without requiring explicit smoothness regularizers or densely sampled sequential data. Extensive experiments demonstrate that TimeFlow outperforms state-of-the-art methods in terms of both future timepoint forecasting and registration accuracy. Moreover, TimeFlow supports novel biological brain aging analyses by differentiating neurodegenerative trajectories from normal aging without requiring segmentation, thereby eliminating the need for labor-intensive annotations and mitigating segmentation inconsistency. TimeFlow offers an accurate, data-efficient, and annotation-free framework for longitudinal analysis of brain aging and chronic diseases, capable of forecasting brain changes beyond the observed study period.
{IEEEkeywords}
Brain MRI Image Registration, Longitudinal Analysis, Brain Aging Analysis
@@ 1 Introduction
Understanding how the human brain evolves throughout the lifespan, whether during healthy aging or in the presence of neuropathological disorders such as neurodegenerative diseases, remains a central challenge in neurology and neuroradiology [1, 2]. Distinguishing these trajectories not only clarifies pathological mechanisms but also refines prognostic models, ultimately supporting timely and personalized interventions [3]. Magnetic Resonance Imaging (MRI) has become the standard tool for capturing longitudinal brain changes, owing to its high spatial resolution and superior soft-tissue contrast. Repeated imaging sessions allow clinicians and researchers to investigate evolving patterns of atrophy, lesion progression, and other structural alterations over time [4, 5].
Figure 1: Top: Given two longitudinal scans (visits at Baseline and Year 4 visits), TimeFlow can interpolate and extrapolate novel brain MRI scans at unseen timepoints (e.g., Year 2 and Year 7). Bottom: Visual comparison of extrapolated predictions at Year 7 demonstrates TimeFlow’s superior performance, compared to (i) NODER, which uses NeuralODE-based geodesic regression, and (ii) VoxelMorph, which linearly scales displacement field. Prediction errors are marked by red arrows.
@@
At the core of most quantitative longitudinal brain analyses lies longitudinal registration, which aligns a subject’s brain scans across multiple timepoints [6, 7]. Despite steady advances in longitudinal registration, two critical challenges persist: 1. Ensuring temporal continuity with limited sequential data. Existing methods often rely on explicit temporal smoothness constraints [8, 9, 10] which usually compromise accuracy. Besides, these methods often require densely sampled longitudinal data (often more than five scans) [11, 12], which is impractical in typical clinical scenarios where usually fewer than three follow-up scans are available. 2. Predicting future brain states and analyzing aging progression. While retrospective assessments of past changes are informative, prospective forecasting of brain anatomy would markedly enhance clinical utility [13]. Such forecasting capabilities enable clinicians to anticipate accelerated disease progression and adjust interventions proactively [14]. However, most recent methods model only the observed scan sequence [15, 16, 17], allowing interpolation between existing timepoints. It limits their capability in extrapolating beyond existing scans, as illustrated in Figure˜1, thereby restricting their ability to assess deviations from expected progression, anticipate disease worsening, and adjust treatment accordingly.
To overcome these limitations, we introduce TimeFlow, a novel learning-based longitudinal brain MRI registration framework. TimeFlow combines a U-Net backbone with temporal conditioning to model neuroanatomy as a continuous function of age. As shown in Figure˜2, TimeFlow requires only two scans per subject to estimate the bidirectional deformation fields for arbitrary timepoints, facilitating both interpolation (timepoints within observed intervals) and extrapolation (timepoints beyond observed intervals). Our proposed bidirectional inter-/extra-polation consistency constraints allow unlimited sampling of timepoints, rather than the limited discrete visits during training, ensuring temporally coherent deformation fields. This synergy between the temporal conditioning and the proposed constraints enables TimeFlow to predict future brain states and supports aging analyses without additional annotations. Unlike generative models that create synthetic MRI scans, TimeFlow uses image warping to ensure anatomically plausible outputs, maintain consistency, and provide interpretability through deformation field inspection.
Our contributions are threefold:
@@
    Temporally-conditioned registration. A unified network conditioned on a continuous time parameter produces temporally coherent deformation fields from only two input scans, accurately predicting unobserved anatomical states.
@@
    Bidirectional inter-/extra-polation consistency loss. Symmetric supervision on bidirectional deformations removes the need for explicit temporal smoothness regularizers, significantly improving temporal coherence, continuity, and generalization through unlimitedly sampled timepoints.
  3. 3.
    Registration-only aging analysis. A registration-only pipeline quantifies relative aging trajectories by referencing healthy controls, differentiating Mild Cognitive Impairment (MCI) and Dementia without requiring tissue labels, thus removing the need for non-trivial annotations.
Figure 2: TimeFlow takes source and target images as input, conditioned on the continuous temporal parameter $t$. Adjusting $t$ enables the generation of interpolated brain scans within the observed interval ($0<t<1$), or extrapolated scans for future ($t>1$).
@@ 2 Related Work
2.1 Registration across subjects
@@
Cross-sectional registration methods laid the groundwork for medical image registration, pioneering this field with techniques such as parametric spline-based models [18, 19], optical flow [20, 21], and diffeomorphic transformations [22, 23, 24]. However, they demand prohibitive computing time in the order of several hours to register one single subject, posing a significant barrier to clinical deployment. To address this, learning-based frameworks [25, 26, 27, 28, 29, 30] have emerged, dramatically reducing inference time and often improving registration accuracy. Nevertheless, all such methods rely on carefully tuned smoothness regularizers to ensure anatomically plausible deformation fields. An improper weighting factor of smoothness regularizers can either reduce fidelity or introduce unrealistic warps. Additional tuning or adaptive strategies are often needed to strike an optimal balance [31, 32, 33].
2.2 Registration across time
@@
Longitudinal registration methods [34, 35, 36, 37, 38, 12, 10, 39] often leverage the backbone architecture of the cross-sectional registration, whereas one additional constraint is introduced: the temporal continuity/consistency across multiple timepoints of a subject. Two major strategies are often used to tackle this problem. First, an explicit smoothness constraint along the temporal dimension is incorporated to maintain coherent deformation fields over time [9, 11]. However, as with spatial smoothness, temporal regularization faces a similar trade-off: enforcing temporal continuity can inadvertently lower spatial accuracy. Second, many longitudinal methods require multiple, densely sampled scans (>10) to capture temporal dynamics [16, 15]. In real-world clinical scenarios, patients often have only a few follow-up scans ($\leq 3$), restricting model flexibility and practical implementation. Importantly, these methods commonly can only carry out retrospective analyses and are not capable of prospective studies, such as future image extrapolation and aging/disease progression risk prediction.
2.3 Longitudinal brain analysis
Existing longitudinal studies of neurodegenerative diseases like Alzheimer’s disease usually require additional annotations, such as cortical surfaces and structural segmentations to quantify the brain changes. However, these labels are usually created by cross-sectional pipelines that process each visit independently, which are both time-consuming and prone to temporal inconsistency, i.e., the same anatomy receives different labels at successive visits. Consequently, quantities derived from those labels (e.g., hippocampal volume, cortical-thickness trajectories) require post-processing to suppress the systematic noise [40, 41]. These limitations motivate annotation-free, temporally coherent models that treat the sequential series as a single spatiotemporal object instead of a collection of disconnected timepoints.
Figure 3: Left: Architecture of TimeFlow. The network $f_{\theta}$ incorporates temporal conditioning via adaptive instance normalization. It takes two input images ($I_{0}$$I_{L}$) and a continuous time variable $t$ to predict the temporal deformation field $\phi_{t}^{0\rightarrow L}$. Examples of $t=0.5$ (interpolation at middle), $t=1$ (endpoint registration), and $t=2$ (extrapolation) are shown. Right: Interpolation (upper yellow) and Extrapolation (bottom green) Consistency Constraints. The symmetric intermediate image similarity enforces the deformed source and target to be aligned at time $t_{k}$, while the interpolation flow consistency loss encourages the deformation composition consistency. The extrapolation similarity and flow constraints give information on what a realistic and reasonable forecast should be. Training with both consistency losses enables TimeFlow to generate plausible deformation in $t\in[-\infty,+\infty]$. The input images are highlighted by blue squares (Input), while the supervision is highlighted by dashed squares (Sup.).
2.4 Future prediction via generative models
Recent advances in diffusion-based generative models have inspired several attempts to synthesize longitudinal brain MRIs conditioned on variables such as age and diagnosis [42, 43]. Despite producing visually plausible synthetic images, these methods largely remain black-box, as their diffusion trajectory does not map clearly onto biophysical deformations. This limits the interpretability, particularly for clinicians interested in tracking local changes like region-specific atrophy. Besides, diffusion models typically require large-scale datasets for stable training, which is impractical in clinical settings where follow-up is sparse. Nonetheless, generative models retain promise as complementary tools in neuropathological disorders involving intensity anomalies, such as gliomas or lesions, where longitudinal registration alone is insufficient to represent newly emerging tissue structures.
@@ 3.1 Time-Continuous Registration
Given a longitudinal MRI sequence $\{I_{k}\mid k=0,\ldots,L\}$, our goal is to derive a temporally continuous deformation trajectory connecting all intermediate timepoints. To achieve this, we propose a time-continuous registration framework, TimeFlow, defined as $\phi_{t}=f_{\theta}(I_{0},I_{L},t)$. As illustrated in Figure˜2, TimeFlow accepts the baseline image $I_{0}$ (source) and the last image $I_{L}$ (target) and predicts the deformation fields $\phi_{t}$ parametrized by continuous time $t$. For interpolation within the observed interval ($0<t<1$), TimeFlow predicts intermediate states between source and target scans. For extrapolation ($t>1$), TimeFlow predicts future brain states beyond the last observed visit. To enforce temporal coherence and continuity in the deformation trajectories, we introduce the novel inter-/extra-polation consistency constraints. The core architecture (Figure˜3, Left) is a time-embedded U-Net [44, 45]. Specifically, a sinusoidal positional encoding and multi-layer perceptrons (MLPs) project the scalar time variable $t$ into a latent embedding space. Next, the time latent vector is fed to two separate MLPs to steer the mean and variance of the inter-layer adaptive instance normalization. The network is thus instilled with the capability of generating meaningful time embedding and predicting time-continuous flow fields by the novel time conditioning mechanism.
3.2 Interpolation Consistency
Symmetric Intermediate Similarity
We first define the time-continuous registration function as $\phi_{t}^{0\rightarrow L}=f_{\theta}(I_{0},I_{L},t)$, where $f_{\theta}(I_{0},I_{L},0)=\textrm{Id}$ is the identity mapping at $t=0$. When $t\in(0,1]$, $\phi_{t}^{0\rightarrow L}$ is the forward deformation mapping from $I_{0}$ towards $I_{L}$, and $\phi_{1}^{0\rightarrow L}$ moves $I_{0}$ to reach $I_{L}$. Similarly, for negative values, $t<0$, $\phi_{t}$ maps backward from $I_{L}$ to $I_{0}$. For any intermediate image $I_{k}$ observed at time $t_{k}$, we can formulate the symmetric interpolation image similarity measure as:
@@ 3.2 Interpolation Consistency
which deforms both source and target images to remove the bias introduced by single-side warping [46, 47, 48]. As shown in Figure˜3 (Right, top yellow), $I_{0}\circ\phi_{t_{k}}^{0\rightarrow L}$ is warping $I_{0}$ to the time $t=t_{k}$ and $I_{L}\circ\phi_{t_{k}-1}^{0\rightarrow L}$ is warping $I_{L}$ to the time $t=t_{k}$. This symmetry enforces bidirectional consistency, facilitating plausible deformation trajectories passing through observed intermediate scans.
Generalized Symmetric Intermediate Similarity
However, in real-world clinical scenarios, many subjects have fewer than three MRI scans available. In such a case, we generalize the above constraint by synthetically sampling intermediate timepoints $\hat{t}$ from a uniform distribution $\mathcal{U}(0,1)$, defining the generalized interpolation similarity as:
@@
Here, both warped images $I_{0}\circ\phi_{\hat{t}}^{0\rightarrow L}$ and $I_{L}\circ\phi_{\hat{t}-1}^{0\rightarrow L}$ serve as a synthetic version of intermediate timpoint $I_{\hat{t}}$, eliminating reliance on actual intermediate observations. By freely sampling $\hat{t}$, TimeFlow effectively learns densely and continuously across the entire temporal axis.
Interpolation Flow Consistency
To further encourage plausible time-continuous deformations, we propose the interpolation flow consistency as a regularizer:
which combines the two constraints shown in Figure˜3 (Right, top yellow). We formulated the loss with the notation $t_{k}$, but as described in the previous paragraph, $t_{k}$ can also be sampled from the uniform distribution if no intermediate observation exists. It is derived from the common transformation composition consistency that two consecutive transforms are equivalent to a direct transform. Moving from $t=0$ to $t=1$ and then back to $t=t_{k}$ is equivalent to moving directly from $t=0$ to $t=t_{k}$ (arrow 1 composed with arrow 3 should equal to arrow 2 in Figure˜3, Right). The backward case is derived similarly. It is worth noting that this differs from the commonly used smoothness regularizations, which explicitly enforce a smooth gradient over the flow field. Therefore, there is only a minimal trade-off between accuracy and smoothness since both the symmetric image similarity and the flow consistency encourage plausible and reasonable longitudinal transformations.
3.3 Extrapolation Consistency
Through the intermediate similarity and interpolation flow consistency, we obtain meaningful predictions for bidirectional registrations in $t\in[-1,1]$. Now we introduce extrapolation constraints for forward ($t>1$) and backward (t<1t<-1) directions. To this end, we consider the deformation $\phi_{1}^{0\rightarrow k}$ from $I_{0}$ to $I_{k}$. Given the time of $I_{k}$ being $t_{k}$, it holds that $\phi_{t_{k}}^{0\rightarrow L}=\phi_{1}^{0\rightarrow k}$. As illustrated in Figure˜3 (Right), they both register $I_{0}$ to $I_{k}$. Consequently, we can then formulate $\phi_{1}^{0\rightarrow L}=\phi_{1/t_{k}}^{0\rightarrow k}$ (arrow 1 equals to arrow 5 in Figure˜3, Right), where $\phi_{1/t_{k}}^{0\rightarrow k}$ denotes an extrapolation because $t_{k}\in(0,1)$ and $1/t_{k}>1$. Based on this notation, we define the extrapolation image similarity as:
Again, if no intermediate observations are available, we can sample $t_{k}$ from $\mathcal{U}(0,1)$ and define $I_{k}=I_{0}\circ\phi_{t_{k}}^{0\rightarrow L}$ to create synthetic intermediate data. In addition to the extrapolation similarity measure, we define the forward extrapolation flow consistency as:
As shown in Figure˜3 (Right), both $\phi_{1}^{0\rightarrow L}$ (arrow 1) and $\phi_{1/t_{k}}^{0\rightarrow k}$ (arrow 5) map from $I_{0}$ to $I_{L}$, thus should be consistent with each other. The backward versions of the extrapolation similarity and extrapolation flow consistency are similarly defined as:
The similarity measures and flow consistencies both contribute to achieving a realistic and coherent extrapolation forecasting. The proposed losses effectively ensure accurate extrapolation while implicitly capturing temporal consistency and continuity, without the need for additional temporal smoothness constraints.
3.4 Unlimited Sampled Triplets
As illustrated in Figure˜3 (Right), the proposed inter-/extra-polation constraints require observed triplets of timepoints, limiting learning when few discrete scans are available. To address this, we use “synthetic” triplets dynamically sampled from continuous time intervals. Specifically, by sampling intermediate timepoints $\hat{t}$ from $\mathcal{U}(0,1)$ during training, we can generate infinite triplets from a pair of source and target images, freeing TimeFlow from the discrete, fixed timepoints. This continuous, densely sampled temporal supervision significantly enhances the network’s capacity to generalize and accurately model brain anatomy changes at any arbitrary timepoint, including unseen future scenarios, while using only two scans as inputs.
4 Data and Experiments
Data and Pre-processing
We train and evaluate TimeFlow on the public Alzheimer’s Disease Neuroimaging Initiative (ADNI) dataset [4], which comprises longitudinal T1-weighted Magnetic Resonance (MR) scans from both healthy subjects and patients diagnosed with Alzheimer’s Disease (AD) or Mild Cognitive Impairment (MCI), i.e., neurodegenerative disorders that exhibit specific patterns of atrophy development in the brain. We employ the longitudinal processing from FreeSurfer v7.2 [49]. First, all scans are processed individually, including skull stripping, normalization. Subsequently, the image sequence is rigidly aligned via a groupwise rigid registration [50]. FreeSurfer aggregates information across time points and refines the pre-processed outputs. To ensure meaningful longitudinal analysis, we select 134 subjects exhibiting substantial structural changes (median brain deformation >0.3>0.3mm). Then, the dataset is divided into training, validation, and testing sets, ensuring a balance in diagnosis, age, sex, and the number of visits. It results in 74, 30, 30 subjects for each split, respectively, with the number of time points per subject ranging from 2 to 12. All scans have isotropic 1mm spacing and are cropped to a uniform size of $160\times 160\times 192$ voxels. The intensity values are scaled to the range [0,1], from percentile [0,99.9] without clipping. Additionally, 11 subjects from the OASIS dataset [51] serve for zero-shot evaluation.
Figure 4: Visualization of extrapolation errors compared to ground truth future images. Coronal and axial views of two subjects are presented. Areas with high error are marked by red arrows.
Implementation Details
Our model is implemented using Pytorch 2.1.2. and training/test is carried out on one NVIDIA A100 40G with a batch size of $2$ (sample $\hat{t}$ twice per subject). We set the dimension of time embedding to $16$, the feature channels in the U-Net encoder and decoder to $[32,32,48,48,96]$. In addition to TimeFlow, which outputs the time-dependent deformation field directly, we have the diffeomorphic variant TimeFlowdiff, which utilizes the scaling and squaring layer with $7$ integration steps to integrate the temporal stationary velocity field. The weights of both interpolation and extrapolation similarity are set to $1.0$. The weights of interpolation and extrapolation flow consistency are set to $2,0.03$ for TimeFlow and $1.25,0.025$ for TimeFlowdiff.
Figure 5: Future image predictions at different extrapolation time $t$, where no ground truth is available. $t=1$ represents one-step pairwise registration between $I_{m}$ and $I_{f}$. Thus, $t=2,4,6$ represents two-, four-, and six-step extrapolation over $I_{f}$. Major errors are highlighted with red arrows. VXM/diff produces unsmooth deformations, leading to structural inconsistencies and loss of anatomical details (zoom in for more details).
[…]